# Efficient Algorithm Level Error Detection for Number-Theoretic Transform Assessed on FPGAs

### v1

@@ III-1 Component-wise NTT Multiplication
We note that (4) is used in our proposed error detection scheme.
If we rotate the input sequence of (1) by one, then we get:
$\left[\begin{array}[]{c}\widetilde{a}(0)\\ \widetilde{a}(1)\\ \widetilde{a}(2)\\ .\\ \widetilde{a}(N-1)\end{array}\right]$ $=\rho*\left[\begin{array}[]{c}\hat{a}(1)\\ \hat{a}(2)\\ .\\ \hat{a}(N-1)\\ \hat{a}(0)\end{array}\right]$ (15)
or
$\left[\begin{array}[]{c}\widetilde{a}(0)/\omega^{0}\\ \widetilde{a}(1)/\omega^{1}\\ \widetilde{a}(2)/\omega^{2}\\ .\\ \widetilde{a}(n-1)/\omega^{(n-1)}\end{array}\right]$ $=\theta*\left[\begin{array}[]{c}\hat{a}(1)\\ \hat{a}(2)\\ .\\ \hat{a}(n-1)\\ \hat{a}(0)\end{array}\right]$ (26)
@@
where the symbol * denotes matrix multiplication, $\rho=\left[\begin{array}[]{cccc}\omega^{0}&\omega^{0}&...&\omega^{0}\\ \omega^{1}&\omega^{2}&...&\omega^{0}\\ \omega^{2}&\omega^{4}&...&\omega^{0}\\ .&.&...&.\\ \omega^{n-1}&\omega^{2(n-1)}&...&\omega^{0}\end{array}\right]$ in (5), and $\theta=$$\left[\begin{array}[]{cccc}\omega^{0}&\omega^{0}&...&\omega^{0}\\ \omega^{0}&\omega^{1}&...&\omega^{n-1}\\ \omega^{0}&\omega^{2}&...&\omega^{2(n-1)}\\ .&.&...&.\\ \omega^{0}&\omega^{n-1}&...&\omega^{(n-1)^{2}}\end{array}\right]$ in (6). From $\alpha$ (in (1)) + $\beta$ (in (6)), we get the following:
@@
$\left[\begin{array}[]{c}\widetilde{a}(0)(\alpha+\beta\omega^{-0})\\ \widetilde{a}(1)(\alpha+\beta\omega^{-1})\\ \widetilde{a}(2)(\alpha+\beta\omega^{-2})\\ .\\ \widetilde{a}(n-1)(\alpha+\beta\omega^{-(n-1)})\end{array}\right]$ $=\theta*\left[\begin{array}[]{c}\alpha\hat{a}(0)+\beta\hat{a}(1)\\ \alpha\hat{a}(1)+\beta\hat{a}(2)\\ .\\ \alpha\hat{a}(n-2)+\beta\hat{a}(n-1)\\ \alpha\hat{a}(n-1)+\beta\hat{a}(0)\end{array}\right]$ (37)
@@
where $\alpha$ and $\beta$ in (7) are scalars which can be selected arbitrarily. Without losing generality and for the sake of simplicity, we select $\alpha=2$ and $\beta=1$. Let us denote $\hat{a}^{m}$ and $\hat{b}^{m}$, respectively, as the inputs which are shifted by $m$. In other words, we can achieve:
$\widetilde{a}=\frac{1}{(\alpha+\beta\omega^{-k})}NTT(\alpha\hat{a}+\beta\hat{a}^{1})$ (38)
$\widetilde{b}=\frac{1}{(\alpha+\beta\omega^{-k})}NTT(\alpha\hat{b}+\beta\hat{b}^{1})$ (39)
Let us denote $c$ as the component-wise NTT multiplication of $\hat{a}$ and $\hat{b}$. We can get:
$c$ $=NTT(\hat{a})\,\Circle\,NTT(\hat{b})=$
$\frac{1}{(\alpha+\beta\omega^{-k})^{2}}NTT(\alpha\hat{a}+\beta\hat{a}^{1})\,\Circle\,NTT(\alpha\hat{b}+\beta\hat{b}^{1})$ (40)
@@
The proposed error detection scheme is depicted in Fig. 2. The Decoder_2 sub-block multiplies its input with $\frac{1}{(\alpha+\beta\omega^{-k})^{2}}$, which is the inverse of $(\alpha+\beta\omega^{-k})^{2}$ in $\mathbb{R}=\mathbb{Z}[X]/(X^{n}+1)$. The scheme for error detection involves comparing $c(0)$, obtained from (10), with $\sum_{j=0}^{n-1}a(j)\sum_{j=0}^{n-1}b(j)$.
@@
Figure 2: Proposed algorithm level error detection scheme for NTT multiplication module.
III-2 Pre-processor
@@
The pre-processor module performs element-by-element multiplication on two input arrays. In proposing the error detection scheme, we have applied an additional recomputing with shifted operands. As shown in Fig. 3, the encoding and decoding modules constitute shifting which is free in hardware. The pre-processing module considers a slight overhead linked to NTT multiplication. As a result, the error detection scheme employing recomputation does not notably impact the overall performance.
@@
Figure 3: Proposed error detection scheme for pre-process module through recompuation with shifted operands.
IV Simulations and Fault Model
In this section, we present the utilized fault model and the simulation results acquired by implementing Python code to show error detection coverage, employing the presented fault model, and utilizing our efficient error detection schemes in this paper.
IV-A Fault Model
@@
We assume that faults occur in either the complex multiplier or adder, or a specific group of input or output lines in the butterfly module which is used in NTT. As depicted in Fig. 4, faults can happen at modules $\{1,2,3\}$ marked in the figure, which may result in erroneous outputs of a butterfly sub-block used in NTT module. Fault occurrence follow a normal distribution within butterfly modules in Fig. 2. Furthermore, as we provide fault detection schemes over NTT multiplication, faults can happen during component wise multiplication module.
@@
Figure 4: The utilized fault model in this work for the butterfly sub-block.
IV-B Error Coverage Simulation Results
@@
To assess the error coverage, we employed Python3 to perform simulations, applying the described fault models and error detection techniques to both the NTT and pre-process modules. The simulations encompassed a million samples, using parameters identical to the standard Kyber $(n=256)$ and BLISS signature [26] $(n=512)$. As shown in Table I, with increase in the number of faults, we can attain an error detection ratio close to 100% for both modules. Higher error coverage is achievable for identifying burst errors. We have done extensive experiments injecting burst adjacent faults, the results of which are not presented for the sake of brevity.
@@
Table I: Achieved error detection ratios of the proposed schemes with 1, 2, 4, 8, and 16 faults occurrence for one million samples
Parameters ${}^{1}n=256,$${}^{2}\omega=17$, $n=512,\omega=2,401$,
${}^{3}q=3,329$ $q=12,289$
Component Pre-process NTT Pre-process NTT
Mult Mult
1 99.6% 51% 99.7% 53%
Number of 2 99.9% 67% 99.9% 70.6%
faults 4 100% 90.3% 100% 90.9%
8 100% 99% 100% 99%
16 100% 99.9% 100% 99.9%
${}^{1}$$n=$ Polynomial size, ${}^{2}$$\omega=$ Twiddle factor, ${}^{3}$$q=$ Prime number
V Xilinx/AMD FPGA Implementation results

### v2

@@ 3.1.1. Component-wise NTT Multiplication
We note that (7) is used in our proposed error detection scheme. If we rotate the input sequence of (1) by one, then we get:
@@
$\left[\begin{array}[]{c}\hat{f}(0)\\ \hat{f}(1)\\ \hat{f}(2)\\ .\\ \hat{f}(N-1)\end{array}\right]=\rho*\left[\begin{array}[]{c}\tilde{f}(1)\\ \tilde{f}(2)\\ .\\ \tilde{f}(N-1)\\ \tilde{f}(0)\end{array}\right]$or $\left[\begin{array}[]{c}\hat{f}(0)/\omega^{0}\\ \hat{f}(1)/\omega^{1}\\ \hat{f}(2)/\omega^{2}\\ .\\ \hat{f}(n-1)/\omega^{(n-1)}\end{array}\right]=\theta*\left[\begin{array}[]{c}\tilde{f}(1)\\ \tilde{f}(2)\\ .\\ \tilde{f}(n-1)\\ \tilde{f}(0)\end{array}\right]$ where the symbol * denotes matrix multiplication, $\rho=\left[\begin{array}[]{cccc}\omega^{0}&\omega^{0}&...&\omega^{0}\\ \omega^{1}&\omega^{2}&...&\omega^{0}\\ \omega^{2}&\omega^{4}&...&\omega^{0}\\ .&.&...&.\\ \omega^{n-1}&\omega^{2(n-1)}&...&\omega^{0}\end{array}\right]$ and $\theta=$$\left[\begin{array}[]{cccc}\omega^{0}&\omega^{0}&...&\omega^{0}\\ \omega^{0}&\omega^{1}&...&\omega^{n-1}\\ \omega^{0}&\omega^{2}&...&\omega^{2(n-1)}\\ .&.&...&.\\ \omega^{0}&\omega^{n-1}&...&\omega^{(n-1)^{2}}\end{array}\right]$.
@@
From $\theta*\left[\begin{array}[]{c}\alpha\tilde{f}(0)\\ \alpha\tilde{f}(1)\\ \alpha\tilde{f}(2)\\ .\\ \tilde{\alpha f}(n-1)\end{array}\right]$$\,+\,\theta*\left[\begin{array}[]{c}\tilde{\beta f}(1)\\ \beta\tilde{f}(2)\\ .\\ \tilde{\beta f}(n-1)\\ \tilde{\beta f}(0)\end{array}\right]$ , we get the following:
(18) $\left[\begin{array}[]{c}\hat{f}(0)(\alpha+\beta\omega^{-0})\\ \hat{f}(1)(\alpha+\beta\omega^{-1})\\ \hat{f}(2)(\alpha+\beta\omega^{-2})\\ .\\ \hat{f}(n-1)(\alpha+\beta\omega^{-(n-1)})\end{array}\right]$ $=\theta*\left[\begin{array}[]{c}\alpha\tilde{f}(0)+\beta\tilde{f}(1)\\ \alpha\tilde{f}(1)+\beta\tilde{f}(2)\\ .\\ \alpha\tilde{f}(n-2)+\beta\tilde{f}(n-1)\\ \alpha\tilde{f}(n-1)+\beta\tilde{f}(0)\end{array}\right],$
@@
where $\alpha$ and $\beta$ in (8) are scalars which can be selected arbitrarily. Let us denote $\tilde{f}^{m}$ and $\tilde{g}^{m}$, respectively, as the inputs which are shifted by $m$. In other words, we can achieve:
(19) $\hat{f}=\frac{1}{(\alpha+\beta\omega^{-k})}NTT(\alpha\tilde{f}+\beta\tilde{f}^{1}),$
(20) $\hat{g}=\frac{1}{(\alpha+\beta\omega^{-k})}NTT(\alpha\tilde{g}+\beta\tilde{g}^{1}).$
Let us denote $h$ as the component-wise NTT multiplication of $\tilde{f}$ and $\tilde{g}$. We can get:
(21) $h$ $=NTT(\tilde{f})\,\Circle\,NTT(\tilde{g})=\frac{1}{(\alpha+\beta\omega^{-k})^{2}}NTT(\alpha\tilde{f}+\beta\tilde{f}^{1})\,\Circle\,NTT(\alpha\tilde{g}+\beta\tilde{g}^{1}).$
@@
The proposed error detection scheme is depicted in Fig. 2. The Decoder_2 sub-block multiplies its input with $\frac{1}{(\alpha+\beta\omega^{-k})^{2}}$, which is the inverse of $(\alpha+\beta\omega^{-k})^{2}$ in $\mathbb{R}=\mathbb{Z}[X]/(X^{n}+1)$. The scheme for error detection involves comparing $h(0)$, obtained from (11), with $\sum_{j=0}^{n-1}\tilde{f}(j)\sum_{j=0}^{n-1}\tilde{g}(j)$ obtained from (7).
@@
Figure 2. Proposed algorithm level error detection scheme for the NTT multiplication module using Negative Wrapped Convolution.
3.1.2. Pre-processor
@@
The pre-processor module performs element-by-element multiplication on two input arrays. In proposing the error detection scheme, we have applied an additional recomputing with shifted operands. As shown in Fig. 3, the encoding and decoding modules constitute shifting which is free in hardware. The pre-processing module considers a slight overhead linked to the NTT multiplication. As a result, the error detection scheme employing recomputation does not notably impact the overall performance.
3.2. Kyber
To apply the Negative Wrapped Convolution method, it is necessary that $2N\mid(q-1)$. For the Kyber parameters with the prime $q=3329$ and $N=256$, the base field $\mathbb{Z}_{q}$ includes 256-th roots of unity but lacks 512-th roots. Therefore, Negative Wrapped Convolution cannot be used in Kyber. We proposed below our error detection scheme on the Kyber NTT module (equations (2) and (3)).
3.2.1. NTT Module
By expressing equations (2) and (3) as a matrix multiplication, we obtain the following relation where $N=256$ with respect to Kyber parameters:
$f_{even}^{{}^{\prime}}=\left[\begin{array}[]{c}\hat{f}(0)\\ \hat{f}(2)\\ .\\ \hat{f}(252)\\ \hat{f}(N-2)\end{array}\right]=\partial*f_{even}$, and $f_{odd}^{{}^{\prime}}=\left[\begin{array}[]{c}\hat{f}(1)\\ \hat{f}(3)\\ .\\ \hat{f}(253)\\ \hat{f}(N-1)\end{array}\right]=\partial*f_{odd}$, where $f_{even}=\left[\begin{array}[]{c}f(0)\\ f(2)\\ .\\ f(252))\\ f(N-2)\end{array}\right]$,
$f_{odd}=\left[\begin{array}[]{c}f(1)\\ f(3)\\ .\\ f(253)\\ f(N-1)\end{array}\right]$, and $\partial=\left[\begin{array}[]{ccccc}1&\omega&\omega^{2}&...&\omega^{127}\\ 1&\omega^{2b(1)+1}&\omega^{2(2b(1)+1)}&...&\omega^{127(2b(1)+1)}\\ 1&\omega^{2b(2)+1}&\omega^{2(2b(2)+1)}&...&\omega^{127(2b(2)+1)}\\ .&.&.&...&.\\ 1&\omega^{2b(127)+1}&\omega^{2(2b(127)+1)}&...&\omega^{127(2b(127)+1)}\end{array}\right]$.
@@ 3.2.1. NTT Module
Figure 3. Proposed error detection scheme for pre-process module through recompuation with shifted operands.
If we rotate the input sequence of $f_{even}^{{}^{\prime}}$ and $f_{odd}^{{}^{\prime}}$ by one, by using the symmetric and periodicity properties ($\omega^{k+N/2}=-\omega^{k}$ and $\omega^{k+N}=\omega^{k})$ of NTT, we get:
(22) $f_{even}^{{}^{\prime\prime}}=\left[\begin{array}[]{c}\frac{\hat{f}(0)-2f(0)}{\omega^{2b(0)+1}}\\ \frac{\hat{f}(2)-2f(0)}{\omega^{2b(1)+1}}\\ .\\ \frac{\hat{f}(252)-2f(0)}{\omega^{2b(126)+1}}\\ \frac{\hat{f}(N-2)-2f(0)}{\omega^{2b(\frac{N-2}{2})+1}}\end{array}\right]=\partial*\left[\begin{array}[]{c}f(2)\\ f(4)\\ .\\ f(N-2)\\ f(0)\end{array}\right],$
(23) $f_{odd}^{{}^{\prime\prime}}=\left[\begin{array}[]{c}\frac{\hat{f}(1)-2f(1)}{\omega^{2b(0)+1}}\\ \frac{\hat{f}(3)-2f(1)}{\omega^{2b(1)+1}}\\ .\\ \frac{\hat{f}(253)-2f(1)}{\omega^{2b(126)+1}}\\ \frac{\hat{f}(N-1)-2f(1)}{\omega^{2b(\frac{N-2}{2})+1}}\end{array}\right]=\partial*\left[\begin{array}[]{c}f(3)\\ f(5)\\ .\\ f(N-1)\\ f(1)\end{array}\right].$
From $\alpha$$f_{even}^{{}^{\prime}}$ + $\beta$$f_{even}^{{}^{\prime\prime}}$ and $\alpha$$f_{odd}^{{}^{\prime}}$ + $\beta$$f_{odd}^{{}^{\prime\prime}}$ , we get the following:
(34) $\left[\begin{array}[]{c}\alpha\hat{f}(0)+\frac{\beta(\hat{f}(0)-2f(0))}{\omega^{2b(0)+1}}\\ \alpha\hat{f}(2)+\frac{\beta(\hat{f}(2)-2f(0))}{\omega^{2b(1)+1}}\\ .\\ \alpha\hat{f}(252)+\frac{\beta(\hat{f}(252)-2f(0))}{\omega^{2b(126)+1}}\\ \alpha\hat{f}(N-2)+\frac{\beta(\hat{f}(N-2)-2f(0))}{\omega^{2b(127)+1}}\end{array}\right]$ $=\partial*\left[\begin{array}[]{c}\alpha f(0)+\beta f(2)\\ \alpha f(2)+\beta f(4)\\ .\\ \alpha f(252)+\beta f(N-2)\\ \alpha f(N-2)+\beta f(0)\end{array}\right],$
(45) $\left[\begin{array}[]{c}\alpha\hat{f}(1)+\frac{\beta(\hat{f}(1)-2f(1))}{\omega^{2b(0)+1}}\\ \alpha\hat{f}(3)+\frac{\beta(\hat{f}(3)-2f(1))}{\omega^{2b(1)+1}}\\ .\\ \alpha\hat{f}(253)+\frac{\beta(\hat{f}(253)-2f(1))}{\omega^{2b(126)+1}}\\ \alpha\hat{f}(N-1)+\frac{\beta(\hat{f}(N-1)-2f(1))}{\omega^{2b(127)+1}}\end{array}\right]$ $=\partial*\left[\begin{array}[]{c}\alpha f(1)+\beta f(3)\\ \alpha f(3)+\beta f(5)\\ .\\ \alpha f(253)+\beta f(N-1)\\ \alpha f(N-1)+\beta f(1)\end{array}\right].$
Let us denote $f_{even}^{1}$ and $f_{odd}^{1}$ , respectively, as the $f_{even}$ and $f_{odd}$ that are shifted by $1$. If we call equations (2) and (3) NTT_Kyber, we can achieve:
(46) $\hat{f}(2k)=\text{$\frac{1}{(\frac{\beta}{\omega^{2b(k)+1}}+\alpha)}$ (NTT\_Kyber}(\alpha f(2k)+\beta f(2k)^{1})+\frac{2\beta f(0)}{\omega^{2b(k)+1}}),\quad k=0,1,...,127,$
(47) $\hat{f}(2k+1)=\text{$\frac{1}{(\frac{\beta}{\omega^{2b(k)+1}}+\alpha)}$ (NTT\_Kyber}(\alpha f(2k+1)+\beta f(2k+1)^{1})+\frac{2\beta f(1)}{\omega^{2b(k)+1}}),\quad k=0,1,...,127.$
We can use the relation presented below for our error detection scheme:
(48) $128(f(0)+f(1))\text{ mod }q=\sum_{j=0}^{n-1}\hat{f}(j)\text{ mod $q$,$\quad q=3329$}.$
The proposed error detection scheme is depicted in Fig. 4. The Decoder_3 sub-block multiplies its input with $\frac{1}{(\frac{\beta}{\omega^{2b(k)+1}}+\alpha)}$, which is the inverse of $(\frac{\beta}{\omega^{2b(k)+1}}+\alpha)$ in $\mathbb{R}=\mathbb{Z}[X]/(X^{n}+1)$. The scheme for error detection involves comparing $128(f(0)+f(1))$ mod $q$, with $\sum_{j=0}^{n-1}\hat{f}(j)$ mod $q$.
Figure 4. Proposed algorithm level error detection scheme for the NTT utilized in Kyber.
4. Error Coverage Simulation Results
4.1. Fault Model
The butterfly module is a key element in the NTT that enables efficient modular transformations, much like its function in the FFT. It processes two input elements by calculating their sum and difference. Its ability to handle modular arithmetic and root of unity multiplications is essential for enhancing the performance of the NTT.
@@ 4.1. Fault Model
As depicted in Fig. 5, faults can happen at modules {1,2, or 3} marked in the figure, which may result in erroneous outputs of a butterfly sub-block used in the NTT module. Fault occurrence follows a normal distribution within butterfly modules in Fig. 2 and Fig. 4. Furthermore, as we provide fault detection schemes over the NTT multiplication, faults can happen during component wise multiplication module. The NTT process involves numerous butterfly operations, where faults may occur at any of these operations. Within each butterfly, faults can happen at positions {1, 2, or 3}. In our fault model, no two faults are allowed to occur in the same butterfly operation. In the normal mode fault injection method, if the NTT process contains $N$ butterfly operations and $F$ represents the total number of faults, each butterfly operation has a probability of $\frac{F}{N}$ of being faulty. Moreover, if a butterfly is faulty, the probability of the fault occurring at position {1, 2, or 3} is $\frac{1}{3}$.
In the burst mode fault injection method, once a faulty butterfly operation is randomly chosen, all subsequent butterfly operations will also be faulty.
@@
Figure 5. The utilized fault model in this work for the butterfly sub-block.
4.2. Fault Simulation
To assess the error coverage, we employed Python3 to perform simulations, applying the described fault models and error detection techniques to both the NTT and pre-process modules.
4.2.1. Negative Wrapped Convolution
@@
The simulations encompassed a million samples, using parameters identical to the standard Kyber Round 1. As shown in Table 1, with increase in the number of faults, we can attain an error detection ratio close to 100% for both modules. Higher error coverage is achievable for identifying burst errors.
The table provides data comparing the effectiveness of the proposed fault injection on two cryptographic components: the Pre-process stage and the NTT Multiplication stage, based on the number of faults injected (1, 2, 4, 8, or 16). For the Pre-process component, for a single fault, we already get high error coverage of 99.7%. As the number of injected faults increases, the ratio quickly reaches 100%. In contrast, the NTT Multiplication stage is more difficult to protect, but eventually we get to close to 100%. With one injected fault, 53% is the detection ratio. However, the success rate increases as more faults are injected: 70.6% for 2 faults, 90.9% for 4, 99% for 8, and 99.9% for 16 faults.
@@
Table 1. Negative Wrapped Convolution achieved error detection ratios of the proposed schemes with 1, 2, 4, 8, and 16 faults occurrence with normal fault injection method for one million samples
Parameters ${}^{1}n=256,^{2}\omega=3,844$,
${}^{3}q=7,681$
Component Pre-process NTT
Multiplication
1 99.7% 53%
Number of 2 99.9% 70.6%
faults 4 100% 90.9%
8 100% 99%
16 100% 99.9%
[…]
4.2.2. Kyber
In the standard Kyber Round 3 NTT implementation, there are 896 butterfly operations, corresponding to the polynomial size of 256. A fault can potentially occur at any of these 896 butterfly operations. The simulations involved one million samples, using the same parameters as in Kyber Round 3.
As indicated in Table 2, increasing the number of faults leads to an error detection ratio approaching 100% in both normal and burst modes. In Kyber’s NTT implementation, the error detection ratios for various fault occurrences were analyzed using one million samples, with parameters set to $n=256$, $\omega=17,$ and $q=3329$. The results show that as the number of injected faults increases, the error detection ratio improves significantly. With just 1 fault, 74.9% of the errors were detected. This detection rate jumps to 93.45% with 2 faults. With 4 faults, the detection ratio reaches 99.49%, and by injecting 8 faults, it further increases to 99.95%. Finally, with 16 faults, the error detection ratio hits 100%, indicating complete detection of faults. This trend highlights the robustness of the error detection mechanism in Kyber’s NTT when subjected to increasing numbers of faults. A higher number of faults leads to near-perfect or perfect error detection, suggesting the scheme becomes more effective at fault identification as fault intensity increases.
The NTT in Kyber was also tested using a burst fault model, with error detection ratios evaluated over one million samples. In the burst fault model, multiple consecutive bits in the data are corrupted simultaneously. The number of burst faults was varied, and the error detection ratios achieved by the proposed fault detection scheme are as follows:
  • With 2 burst faults, the scheme detected errors with a success rate of 93.82%.
  • When 3 burst faults occurred, the detection rate increased to 98.3%.
  • With 4 burst faults, the detection ratio improved further to 99.42%.
  • At 5 burst faults, the scheme achieved a near-perfect detection rate of 99.76%.
  • For 6 burst faults, the detection rate was nearly flawless, reaching 99.8%.
The trend shows that as the number of burst faults increases, the error detection capability of the scheme improves, approaching almost 100% accuracy with 5 or more faults. This indicates the robustness of the detection method, particularly when dealing with higher levels of fault occurrences.
Table 2. NTT in Kyber achieved error detection ratios of the proposed schemes with normal and burst fault injection methods for one million samples
Parameters $n=256,\omega=17$
$q=3,329$
Fault injection method Normal mode Burst mode
1 74.9% 2 93.82%
Number of 2 93.45% 3 98.3%
faults 4 99.49% 4 99.42%
8 99.95% 5 99.76%
16 100% 6 99.8%
5. Xilinx/AMD FPGA and Software Implementation results
5.1. Negative Wrapped Convolution
To confirm the efficiency of our presented approach, we chose to assess its performance by applying it to the NTT multiplication involving a 256-degree polynomial. We conducted a benchmark for implementation on different FPGAs, i.e., Xilinx/AMD Zynq Ultrascale+ and Artix-7. The Xilinx/AMD Zynq Ultrascale+ is a high-performance, multi-core SoC (System on Chip) that integrates programmable logic with ARM Cortex-A53 processors. It is designed for complex applications requiring significant processing power, such as AI, 5G, and automotive systems. It offers advanced features like multi-core processing, high-speed I/O, and power optimization.
In contrast, the Xilinx Artix-7 is a lower-cost, mid-range FPGA aimed at applications that need efficient power usage and moderate performance, such as communications, video processing, and embedded systems. It does not include integrated processors, relying solely on programmable logic.
Key differences include the Zynq Ultrascale+’s integrated ARM processors and higher performance capabilities, whereas the Artix-7 focuses on power efficiency and simpler, more cost-effective designs.